OpenStudy (ms_sp):

1. The sale price of a coffeemaker is $22.75. This is 30% off of the original price. What was the original price of the coffeemaker? 2. One hundred forty countries participated in 1984 Summer Olympics in Los Angeles. This was 75% more than the number of countries that took part in the Summer Olympics in Moscow 4 years earlier. How many countries participated in the 1980 Olympics in Moscow? 3. In 2103, there were 19,767 Starbucks stores world-wide. This is approximately 4551% more stores than in 1994. How many Starbucks stores were there in 1994? (Round to the nearest whole number.)

OpenStudy (ben00):

I think number 1 is easy.

OpenStudy (ben00):

to find the total price have to change 30% to a decimal that would be 0.30

OpenStudy (ben00):

then multiply 22.75 x 0.30 = 6.825 (round it to the nearest cent)

OpenStudy (ben00):

the add 22.75 + 6.83

OpenStudy (ben00):

= $29.58 as the orginal price.

OpenStudy (ben00):

medal me if I am correct.

OpenStudy (ms_sp):

1. 32.50 2. 80 3. 425

OpenStudy (ms_sp):

@alayrhi

OpenStudy (alayrhi):

what do you need help with?

OpenStudy (ms_sp):

The 1-3 questions

OpenStudy (alayrhi):

k the answer to 1 is 32.50.

OpenStudy (alayrhi):

Sorry idk the answers to the others. But i believe @Ben00 is correct.

OpenStudy (ben00):

Then medal me! :D

OpenStudy (ms_sp):

@welshfella

OpenStudy (ms_sp):

@LannyXX or @Kevin HELP

OpenStudy (lannyxx):

Why not try convert each of these statements into algebraic symbols? Convert English to math!! Here this is the first one: salePrice = $22.75 salePrice = 30% "of" the original price = 30% * original find original ???

OpenStudy (lannyxx):

Try to convert each into math, we can check your work!

OpenStudy (ms_sp):

So, 22.75 x .03?

OpenStudy (lannyxx):

30% is 30/100 right? salePrice = 30/100 * original. how do you solve for original? if you need to put it in one side, what happens to 30/100?

OpenStudy (lannyxx):

OK. Let's take this step by step, tell me where exactly you get confused. English Statement to convert to math: "The sale price of a coffeemaker is $22.75. This is 30% off of the original price. What was the original price of the coffeemaker?" 1. Identify your variables. a) sale price b) original price sale price is known, original price is unknown. you need to find it! :D 2. convert "The sale price of a coffeemaker is $22.75 this tells us: salePrice = $22.75 "THIS (sale price) is 30% off of the original" ohhhh. salePrice = 70% original see, it's 30% OFF, so you get the saleprice for 70% the price!!! Do you get it so far?

OpenStudy (lannyxx):

so you go to buy something, and they're giving you 30% off. 100% - 30% = 70%. They're selling their product at 70% the original price

OpenStudy (ms_sp):

Here's what I have so far: 22.75 x 100/ 70

OpenStudy (lannyxx):

yeah!!! salePrice = 70/100 * originalPrice you multiply both sides by 100/70! salePrice * 100/70 = originalPrice you're right, that's the answer for 1.

OpenStudy (lannyxx):

try to follow the procedure for 2, see if you can identify the knowns and unkowns, then see if you can make an mathematical expression based on them!

OpenStudy (lannyxx):

it's fine!

OpenStudy (ms_sp):

Haha thanks, so was I right before I messed up?

OpenStudy (lannyxx):

2. One hundred forty countries participated in 1984 Summer Olympics in Los Angeles. This was 75% more than the number of countries that took part in the Summer Olympics in Moscow 4 years earlier. How many countries participated in the 1980 Olympics in Moscow? Known quantities: number of participant countries in 1984 = 140 countries Unknown quantities: number countries participating in Moscow 4 years earlier << FIND (Let's call this x) Statements: 140 = 75% * x + x Find x??? 140 = 75% x + 100% x = 175% x x = 140/175%

OpenStudy (lannyxx):

Remember, it said 75% more. Remember the sales? it was 30% less than the original 100% - 30% = 70% This is saying it's 175% as much! or the same, and 75% more!

OpenStudy (lannyxx):

the same trend follows with number 3.... when it says there is 50% more than that for example I have 50% more cookies than you!!! It means the number of cookies I have is the number of cookies you have + half of the number of cookies you have. If I was 100% taller than you. Then If you're like, say, 100 meters. I'd be 200 meters. Cuz 100% meters + 100% meters = 2 * 100 = 200 meters. I hope those hints can help you solve those problems!!

OpenStudy (ms_sp):

Kind of confusing there 's two more problems but I am trying them out currently.

OpenStudy (lannyxx):

Yeah I know. Word problems can be confusing, but with practice you'll become a pro! Maybe if you can try to explain it to me, we could pinpoint where you're confused. then we can work on that :P

OpenStudy (ms_sp):

So what you told me so far was confusing for number 3

OpenStudy (ms_sp):

The way I am trying solve it is not working.

OpenStudy (lannyxx):

why not? 3. In 2103, there were 19,767 Starbucks stores world-wide. This is approximately 4551% more stores than in 1994. How many Starbucks stores were there in 1994? (Round to the nearest whole number.) in 2103: We had 19,767 stores That is 4551% more than we had in 1994.... Let x = numStores in 1994 19767 = x + 4551% x = x ( 1 + 4551/100)

OpenStudy (ms_sp):

Oh, because the answer is 425. And I thought you subtract both sides.

OpenStudy (ms_sp):

@NerdyChick16

OpenStudy (ms_sp):

I am still not getting the correct answer.

OpenStudy (ms_sp):

Oh I got the answer! :D Thank you all! @LannyXX , so what I did was 19767=4551 x+x, 19767 = 4551%+100% =4651%, then I did was 19767/4651 = 425.

OpenStudy (lannyxx):

Hurrah!! Also you meant to say 19767/4651% lol or 19767 * 100 / 4651. Good job~

OpenStudy (ms_sp):

Yes! :P